OpenStudy (anonymous):

Gretchen bought prizes for games at her birthday party. She bought 6 bracelets and 3 hair clips. She spent a total of $13.71. Part 1: Write the equation for this word problem. Be sure to identify your variables. Part 2: If each hair clip cost $0.99, solve the equation to find the cost of each bracelet. Show your work.

OpenStudy (hoblos):

b=price of bracelets c=price of hair clips 6b+3c=13.71 if c=0.99 6b=13.71-2.97 b=$1.788